write an equation of the line in slope-intercept form slope:-1 passes through (4,-7)
Answer by math_tutor2020(3817)


y - y1 = m(x - x1) ... Point Slope form

y - y1 = -1(x - x1) ... plugging in the given slope

y - (-7) = -1(x - 4) ... plugging in the given point

y + 7 = -1(x - 4)

y + 7 = -1*x -1*(-4) ... distribute

y + 7 = -x + 4

y + 7 - 7 = -x + 4 - 7 ... subtracting 7 from both sides

y = -x - 3

Answer: y = -x - 3